# MATCHA: Can Multi-Agent Collaboration Build a Trustworthy Conversational Recommender?

### v1

@@ 6 Results
Table 1: Results for top @5 (above) and @10 (below) recommendations
Table 1 highlights the performance of various methods across multiple metrics for top-5 and top-10 recommendations. MATCHA consistently outperforms other methods in most of metrics, showcasing its robustness in conversational recommendation systems. For factuality, MATCHA achieves near-perfect scores (.99), matching OMuleT and surpassing other baselines, indicating its reliability in generating accurate recommendations. In relevance, MATCHA excels with the highest Hit@5 (.29) and Precision@5 (.10) scores, demonstrating its ability to closely align recommendations with user preferences. Similarly, its novelty scores, RPop50@5 (2.05) and RPop50@10 (1.81), indicate a strong balance between recommending popular and niche games, enhancing user satisfaction with diverse options. For coverage, MATCHA achieves competitive entropy (E@5: 8.40, E@10: 8.65) while minimizing repetition (MaxF@5: .09, MaxF@10: .17). Its jailbreak prevention capabilities () and top-tier explanation scores (Exp@5: 4.2, Exp@10: 4.1) further distinguish it, ensuring safety, transparency, and user trust. Compared to baselines like MACRS-C and MACRec, MATCHA emerges as a comprehensive solution excelling across most key metrics for game recommendation. In coverage, OMuleT slightly surpasses MATCHA in entropy for top-10 recommendations (E@10: 9.21 vs. 8.65), indicating that OMuleT provides a slightly more diverse pool of recommendations at a broader scale. MATCHA also achieves an explanation score of 4.1 as evaluated by a virtual judge. Furthermore, we conducted a human evaluation [^1] See Appendix C for details on the human annotation process and examples., where domain experts assigned an average score of 3.97 to MATCHA’s recommendations and its reason. This result aligns closely with the virtual judge evaluation, although a small gap persists, indicating room for improvement in generating explanations that persuade users to engage with or select the recommended games. This highlights an future area for refinement in enhancing user trust and interaction.
@@ 8 Deployment
Our application is built on a full-stack server using Streamlit, which simplifies the creation of an interactive user interface and the management of backend operations. The deployment is hosted in an internal data center, using HashiCorp’s Nomad and Consul for cluster orchestration, deployment, and configuration management. To enhance usability, we have improved the feedback mechanism, allowing users to provide responses on the quality of recommendations directly within the application. This feedback is integrated into the system’s evaluation pipeline, facilitating continuous improvement of the recommendation framework. The application is accessible through the internal VPN, enabling widespread testing and iteration based on diverse user interactions. An example of the system in action is shown in Figure 5, with additional examples provided in the Appendix  D.
Figure 5: Demonstration of the application interface showcasing recommendations with Reasoning and user feedback integration.
@@ 9 Conclusion& Discussion
This work introduces MATCHA, a multi-agent conversational recommendation framework leveraging LLMs with specialized tools. Experiments show MATCHA delivers accurate, diverse, and user-aligned game recommendations while ensuring safety and transparency. Deployment insights lay the groundwork for broader scaling. Differences between virtual judge assessments and human annotations also highlight the need for refining human preference modeling. Future work will enhance explainability, better capture human preferences, explore other domains, and improve the balance between diversity and relevance in recommendations.
@@ 11 Limitations
Appendix A Details of Candidate Generation Tools
The Candidate Generation stage utilizes over ten specialized tools adopted from eun Yoon et al. (2024) to ensure a robust and diverse recommendation pool. In Table  2, we provide a detailed description of each tool and its function.
@@ Appendix A Details of Candidate Generation Tools
Appendix B Explanation Score Evaluation Details
@@
Appendix C Human Evaluation Details
@@
C.1 Human Annotation Example
@@ C.2 Evaluation Interface Demonstration
Appendix D Demo Examples
In this section, we provide additional examples demonstrating the functionality of the application. Each example highlights different scenarios, such as personalized recommendations, rejection to answer to Jailbreak attempt, and feedback integration. See Figure 7 and  8 for an additional illustration.

### v2

@@ 6 Results
Table 1: Overall performance across top-5 (top half) and top-10 (bottom half) recommendations. Bold highlights indicate the best score in each column. $\uparrow$ denotes higher is better; $\downarrow$ denotes lower is better. Checkmark () indicates successful jailbreak defense. Significance threshold: $p<0.01$.
Table 1 highlights the performance of various methods across multiple metrics for top-5 and top-10 recommendations. Additional studies on the MATCHA framework in movie recommendation settings, as well as extended evaluations on the DAN jailbreak dataset, are provided in Appendix H. In high-cardinality recommendation tasks—such as game or app recommendation, where the candidate space may exceed 10,000 items—seemingly small absolute gains in top-k accuracy (e.g., +0.01 in Hit@5) can correspond to hundreds of more relevant results across large-scale deployments. Prior work, has similarly emphasized the impact of modest improvements in top-k relevance metrics eun Yoon et al. (2024); Kook et al. (2025). MATCHA consistently outperforms other methods in most metrics, showcasing its robustness in conversational recommendation systems. For factuality, MATCHA achieves near-perfect scores (.99), matching OMuleT and surpassing other baselines, indicating its reliability in generating accurate recommendations. In relevance, MATCHA achieves the highest Hit@5 (.29) and Precision@5 (.10), demonstrating strong alignment between its recommendations and user preferences. Its novelty scores (RPop50@5: 2.05, RPop50@10: 1.81) reflect a healthy mix of mainstream and niche content, which is crucial for user engagement in interactive settings. For coverage, MATCHA maintains high diversity (E@5: 8.40, E@10: 8.65) while minimizing repetition (MaxF@5: .09, MaxF@10: .17). Its jailbreak prevention capabilities () and strong explanation quality (Exp@5: 4.2, Exp@10: 4.1) further distinguish it in terms of safety and transparency. These results indicate that MATCHA not only improves relevance but also maintains competitive diversity and factual accuracy. In particular, the simultaneous gains in novelty (lower RPop50) and explanation quality suggest that MATCHA generates recommendations that are both surprising and justifiable, which can enhance user satisfaction and trust in interactive settings. Compared to baselines like MACRS-C and MACRec, MATCHA delivers more consistent improvements across key dimensions. While OMuleT slightly surpasses it in entropy (E@10: 9.21 vs. 8.65), MATCHA achieves a more favorable balance of relevance, novelty, and explanation quality. MATCHA also received an explanation score of 4.1 from a virtual judge. A follow-up human evaluation (See Appendix F) by domain experts yielded an average score of 3.97, showing close agreement. This small gap suggests that while machine-evaluated justifications are reliable, further refinements may help better align generated explanations with human expectations and increase user trust. For computational overhead analysis, please see Appendix D.
@@ 8 Deployment
Our application is built on a full-stack server using Streamlit, which simplifies the creation of an interactive user interface and the management of backend operations. The deployment is hosted in an internal data center, using HashiCorp’s Nomad and Consul for cluster orchestration, deployment, and configuration management. To enhance usability, we have improved the feedback mechanism, allowing users to provide responses on the quality of recommendations directly within the application. This feedback is integrated into the system’s evaluation pipeline, facilitating continuous improvement of the recommendation framework. The application is accessible through the internal VPN, enabling widespread testing and iteration based on diverse user interactions. An example of the system in action is shown in Figure 4, with additional examples provided in the Appendix  G.
Figure 4: Demonstration of the application interface showcasing recommendations with reasons.
@@ 9 Conclusion& Discussion
This work introduces MATCHA, a multi-agent conversational recommendation framework leveraging LLMs in combination with specialized tools. Extensive experiments show that MATCHA delivers accurate, diverse, and user-aligned game recommendations while maintaining strong guarantees for safety and transparency. In addition to proposing the core framework, this paper also provides a detailed analysis of system design, computational cost, and key challenges unique to game recommendation. The system has already been deployed internally, offering valuable feedback on performance, robustness, and usability under real-world setting. Future work will enhance explanation quality, better capture nuanced human preferences, explore additional domains, and improve the trade-off between diversity, relevance, and personalization.
@@ 11 Limitations
Appendix A Zero-Shot Hallucination Study: Games vs. Movies
To evaluate the knowledge coverage of LLMs across domains, we conducted a zero-shot recognition and hallucination test comparing game and movie titles.
Setup.
We sampled 5,000 game titles from the Roblox platform and 5,000 movie titles from the ReDial dataset. For each title $t$, GPT-4o was prompted with: “Describe the game/movie $t$ in one sentence.”
Metrics.
  • Recognition Rate: Whether the output included the correct genre or key characteristics.
  • Hallucination Rate: Whether the output introduced fabricated or factually incorrect details.
Results.
Table 2 presents the recognition and hallucination rates. GPT-4o performs substantially worse on game titles, suggesting a coverage gap that supports our hypothesis that games are underrepresented in pretraining data. RR
Domain RR (%) HallR(%)
Games 76.2 23.8
Movies 95.3 4.7
Table 2: Zero-shot recognition and hallucination performance on 5,000 game titles and 5,000 movie titles using GPT-4o. $R$ denotes Recognition Rate and HallR denotes Hallucination Rate.
Appendix B Safety Risks in Game Conversational Recommendation
We further examined the unique safety risks posed by game-focused conversational recommender systems (CRS), both through empirical comparison and external case studies.
Empirical Comparison.
We evaluated GPT-4o on 1,000 user queries for both game and movie recommendation scenarios. Table 3 summarizes the unsafe content rates.
Domain Unsafe Response Rate (%)
Game CRS 7.8
Movie CRS 1.2
Table 3: Proportion of responses containing violent, discriminatory, or inappropriate content.
The results suggest that game CRS systems have a significantly broader attack surface, increasing the likelihood of unsafe or policy-violating outputs.
Real-World Evidence.
These empirical findings are reinforced by documented safety failures in deployed systems:
AI Dungeon Incident Utah Business (2021): Unmoderated game-style generation using GPT-3 led to disturbing outputs, including illegal content, prompting emergency safeguards.
Roblox Case Study Chawki (2025): Analyzed how recommendation and feed systems on game platforms can unintentionally expose users to toxic or grooming-prone content.
Generative NPC Audit Buongiorno et al. (2024): Found that game-based dialogue agents are particularly prone to explicit, biased, or hallucinated outputs in adversarial, multi-turn interactions.
These observations highlight the need for game-specific safeguards in CRS—such as jailbreak prevention, content filtering, age-aware reasoning, and explainability mechanisms—to mitigate both technical and ethical risks.
Appendix C Details of Candidate Generation Tools
The Candidate Generation stage utilizes over ten specialized tools adopted from eun Yoon et al. (2024) to ensure a robust and diverse recommendation pool. In Table  4, we provide a detailed description of each tool and its function.
@@ Appendix C Details of Candidate Generation Tools
Appendix D Computational Cost of Multi-LLM Collaboration & Reflection
This section provides a detailed analysis of the inference latency, resource usage, and cost associated with the MATCHA framework.
Inference.
All LLM calls are executed asynchronously. The two ranking LLMs run in parallel, and the Reflection Agent is invoked only on the top-8 candidates. The table below summarizes average latency, GPU memory usage, and estimated API costs:
Model Variant Latency / Turn GPU Memory Cost (USD)
MATCHA (full) 1.32 s 11.1 GB 0.00012
w/o reflection 1.11 s 10.4 GB 0.00010
w/o multi-LLM 1.09 s 10.3 GB 0.00009
Table 5: Inference efficiency across MATCHA variants.
The full system adds only 0.21 seconds over the single-agent variant—well below typical user response times in dialogue settings.
Training.
MATCHA is training-free. The framework relies solely on curated tool usage and safety rules, with no additional fine-tuning or training cost beyond baseline LLM inference.
Deployment Efficiency.
To ensure scalability, the system includes adjustable knobs: reflection is limited to a single pass, and the candidate pool is capped at 30 items. A hyperparameter table and deployment guidelines will be provided.
Appendix E Explanation Score Evaluation Details
@@
Appendix F Human Evaluation Details
@@
F.1 Human Annotation Example
@@
F.2 Evaluation Interface Demonstration
Appendix G Demo Examples
In this section, we provide additional examples demonstrating the functionality of the application. Each example highlights different scenarios, such as personalized recommendations, rejection to answer to Jailbreak attempt, and feedback integration. See Figure LABEL:fig:appendix_demo_example and  6 for an additional illustration.
Appendix H Additional Evaluation on Jailbreak Robustness and General Recommendation
Evaluation on Real-World Jailbreak Prompts.
To evaluate the effectiveness of MATCHA’s Risk Control Module against adversarial prompts, we conduct additional experiments on the “Do Anything Now” (DAN) benchmark Vargas and Castells (2011), which contains jailbreak prompts collected from 131 real-world online communities between December 2022 and December 2023. These prompts cover a wide range of adversarial strategies, including prompt injection, roleplay-based exploits, and persistent jailbreak patterns observed in the wild. The dataset is considered highly realistic and challenging due to its human-crafted nature and longevity of effective jailbreaks.
Table 6 reports the defense success rate across multiple models and frameworks. MATCHA achieves a success rate of 94.17%, significantly outperforming baseline systems such as GPT-4o (70.13%) and MACRec (74.83%). These results demonstrate MATCHA’s strong generalization to adversarial prompts beyond synthetic test cases, validating the robustness of its multi-layered Risk Control mechanism.
Table 6: Defense success rate on “Do Anything Now” jailbreak dataset
Model / System Defense Success Rate (%)
LLaMA-405B 43.22
GPT-4o 70.13
MACRS-C 63.05
MACRec 74.83
OMuleT 68.50
MATCHA (ours) 94.17
Generalization to Movie Recommendation (ReDial).
To evaluate the generalizability of MATCHA beyond the gaming domain, we adapt the framework to the ReDial dataset, a widely used benchmark for multi-turn movie recommendation. ReDial consists of annotated dialogues where one user seeks movie suggestions and the other responds, making it suitable for assessing open-domain CRS performance.
MATCHA is applied to ReDial with minimal modification, replacing game-specific APIs with a structured movie database. We report two standard metrics: Recall@10 and HitRate@10 (HR@10), which assess the system’s ability to retrieve and surface relevant items within the top-10 predictions. As shown in Table 7, MATCHA achieves the highest Recall@10 score (0.182) and performs competitively on HR@10 (0.289), closely matching the best-performing baseline (MACRS).
Table 7: Recommendation performance on ReDial dataset
Model / System Recall@10 HR@10
LLaMA-405B 0.143 0.172
GPT-4o 0.171 0.278
MACRS 0.174 0.291
MACRec 0.164 0.235
MATCHA (ours) 0.182 0.289
These results suggest that MATCHA’s modular architecture, originally designed for game recommendation, generalizes effectively to other CRS domains. Without domain-specific retraining, MATCHA maintains strong recommendation performance, supporting its flexibility and extensibility. Combined with its demonstrated safety advantages, MATCHA offers a promising foundation for building robust, trustworthy conversational recommenders across domains. g foundation for building safe and effective CRS across diverse domains.